OpenStudy (anonymous):

graph the equation using the point and slope y+7=1/2(x-5)

OpenStudy (anonymous):

what point do you have to use?

OpenStudy (anonymous):

Thats all I got? thats the whole question

OpenStudy (anonymous):

i can show you the easiest way to graph a quadratic function. with intercepts.

OpenStudy (anonymous):

I have no clue what you just said...lol... if you want I can take a screen pic of what the actual prob

OpenStudy (anonymous):

firstly write the equation under this form: \[y=\frac{ 1 }{ 2 } \times x + \frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

y intercept (when the grapf intersects the oy axis): x=0 => y= 9/2 => the first point is A(0, 9/2)

OpenStudy (anonymous):

x intercept (when the grapf intersects the ox axis): y=0 => \[\frac{ x+9 }{ 2 } = 0 => x = - 9\] the second point is B (-9, 0)

OpenStudy (anonymous):